# Qumode Tensor Networks for False Vacuum Decay in Quantum Field Theory

### v1

@@ III.2.2 Phase transitions
Figure 13: Phase transitions on the qumode lattice for two different parameter settings.
@@
Now let us turn to phase transitions in the system as a whole. In Fig. 13 we show two examples of phase transitions that were produced with the lattice initialised in the random sampling approximation to the QFT false vacuum that was discussed in Sec. II.6. The potential used for these examples is the $\tanh q^{\ell}$ function of Eq. (24) with $\ell=2$. The true minimum for these examples was chosen to be at $\varphi_{\rm tv}=0.5$ in order to allow for relatively rapid transitions. The figures show the production of bubbles which coalesce to complete the phase transition. The first example with $\varepsilon=0.5$ is very rapid because the instanton is very small (with $r_{c}\approx 2$). The second example has $\varepsilon=0.4$ and the instanton size has grown enough ($r_{c}\approx 5$) to noticeably slow the bubble production, as would be anticipated from the analysis in Sec. III.1. One observation that supports the bubble formation being the result of genuine instanton transitions is that we find that the maximum bond dimension has to exceed a certain threshold in order for this behaviour to appear (if the dimension is set to be too small then there is little tunnelling). Moreover, this threshold appears to grow with the instanton size. In future work, an in-depth study of these relations will be carried out, with a particular focus on confirming the exponential dependence of the decay rate $\Gamma$ on the action.
IV Summary and Conclusions
@@
In this work, we have presented a novel framework for simulating false vacuum decay in scalar QFT using a qumode-based Hamiltonian formulation. Our method has the potential to overcome longstanding challenges associated with simulating highly entangled, non-equilibrium tunnelling dynamics in QFT by leveraging a qumode lattice representation and implementing real-time evolution via tensor network techniques, specifically the time-evolving block decimation algorithm.
@@
Unlike previous approaches, which typically restricted attention to spin chains or Euclidean simulations, our framework, based on Hamiltonian simulation, remains firmly within the QFT setting and directly simulates real-time tunnelling phenomena. The core innovation lies in the use of qumodes-bosonic modes with (quasi-)continuous Hilbert spaces, which enable the construction of scalar field theories with arbitrary potentials. This setup can, in principle, be implemented on continuous-variable quantum computing devices in its presented form. However, to enable simulation on current classical hardware, we truncate the local Hilbert space and employ a matrix product state ansatz for the quantum state, thus achieving scalable simulations with tunable entanglement control.
We demonstrated the validity of our method through a sequence of increasingly complex tests. First, we validated the real-time dynamics of isolated qumodes in double-well potentials, including adiabatic preparation of metastable states and analytic tunnelling in exactly solvable potentials such as the Pöschl-Teller well. We then extended the method to full (1+1)-dimensional QFTs on qumode lattices, capturing bubble nucleation and expansion, and reproducing analytic expectations such as hyperbolic wall trajectories and critical bubble dynamics.
@@
Importantly, we showed that quantum tunnelling emerges in this framework only once sufficient entanglement is captured by the MPS. The bond dimension required increases with instanton size, suggesting a direct numerical signature of the entanglement entropy scaling expected for false vacuum decay. This provides a powerful diagnostic and highlights the necessity of entanglement-aware methods in simulating quantum phase transitions.
@@
Compared to previous literature, our work represents the first demonstration of scalar QFT tunnelling using a real-time evolved Hamiltonian lattice system with quasi-continuous local degrees of freedom. While continuous-variable tensor network methods, such as continuous MPS or cPEPS [80, 81], have been developed for ground-state field theory, and continuous-variable quantum computing proposals exist for QFT [24, 82], no prior work has combined qumode-lattice discretisation with real-time tensor-network evolution to simulate false vacuum decay in QFT. Our method is thus, to the best of our knowledge, the first demonstration of qumode tensor networks applied to non-equilibrium tunnelling dynamics in scalar field theory.
It can open the door to systematic studies of vacuum decay in arbitrary scalar potentials, including non-renormalisable theories and those beyond the reach of semiclassical methods. Moreover, our framework is designed to be portable to CVQC platforms, where, if technically realised at scale and high fidelity, genuine continuous-variable computing would be able to improve on tensor network simulations in scenarios with large entanglement.
@@
Future extensions include implementing entangled initial state preparation via full-lattice adiabatic evolution, exploring multi-field tunnelling landscapes, and applying this method to higher-dimensional QFTs. Our results show the viability of qumode tensor networks as a bridge between classical simulation and future quantum computing applications in non-perturbative quantum field theory.
@@ A.1 Free fields on the qumode lattice
$\mathscr{L}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}\left(\partial^{\mu}\varphi(x,t)\partial_{\mu}\varphi(x,t)-\omega^{2}\varphi(x,t)^{2}\right)\leavevmode\nobreak\ ,$ (57)
@@
$\pi(x,t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\partial\mathscr{L}}{\partial(\partial_{0}\varphi(x,t))}\leavevmode\nobreak\ =\leavevmode\nobreak\ \partial_{0}\varphi(x,t)\leavevmode\nobreak\ ,$ (58)
@@
$\mathscr{H}(x,t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}\left(\pi(x,t)^{2}+(\nabla\varphi(x,t))^{2}+\omega^{2}\varphi(x,t)^{2}\right)\leavevmode\nobreak\ .$ (59)
@@
$\hat{q}_{n}(t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \varphi(x_{n},t)\leavevmode\nobreak\ ,\qquad\hat{p}_{n}(t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \pi(x_{n},t)\leavevmode\nobreak\ .$ (60)
@@
$\partial_{i}\varphi(x_{n},t)\leavevmode\nobreak\ \leftrightarrow\leavevmode\nobreak\ \frac{{\hat{q}}_{n+1}(t)-{\hat{q}}_{n}(t)}{a}\leavevmode\nobreak\ .$ (61)
@@
$H\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum_{n=0}^{N-1}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,\overline{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+\omega^{2}\hat{q}_{n}(t)^{2}\right]\leavevmode\nobreak\ ,$ (62)
@@
$[\hat{q}_{n},\,\hat{p}_{m}]\leavevmode\nobreak\ =\leavevmode\nobreak\ i\delta_{nm}\leavevmode\nobreak\ ,$ (63)
@@
$\overline{n+1}\leavevmode\nobreak\ =\leavevmode\nobreak\ n+1\mod(N)\leavevmode\nobreak\ .$ (64)
@@
$|\Psi\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \otimes_{n}|\psi_{n}\rangle\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |{\mathbf{q}}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \otimes_{n}|q_{n}\rangle\leavevmode\nobreak\ .$ (65)
@@
[…]
@@
$\langle{\mathbf{q}}|\Psi_{\rm degen}\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n}\langle q_{n}|\psi_{n}\rangle\leavevmode\nobreak\ .$ (67)
@@
$U_{\alpha m}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\sqrt{N}}\nu^{\alpha m}\leavevmode\nobreak\ ,$ (68)
@@
$H\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{a}{2}\sum^{N-1}_{\alpha=0}\left(|\hat{p}_{\alpha}|^{2}+\omega_{\alpha}^{2}|\hat{q}_{\alpha}|^{2}\right)\leavevmode\nobreak\ ,\$ (69)
@@
$\hat{p}_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{n}U_{\alpha n}\hat{p}_{n}\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \hat{q}_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{n}U_{\alpha n}\hat{q}_{n}\leavevmode\nobreak\ ,$ (70)
@@
$\omega_{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\left(\omega^{2}+\frac{4}{a^{2}}\sin^{2}\left(\frac{\pi\alpha}{N}\right)\right)}\leavevmode\nobreak\ .$ (71)
@@
[…]
@@
$\hat{a}_{n}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}+\frac{i}{\omega}\hat{p}_{n}\right)\leavevmode\nobreak\ ,$
$\hat{a}^{\dagger}_{n}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}-\frac{i}{\omega}\hat{p}_{n}\right)\leavevmode\nobreak\ ,$ (73)
@@
$\hat{a}_{\alpha}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega_{\alpha}}{2}}\leavevmode\nobreak\ \sum_{n}\left(U_{\alpha n}\,\hat{q}_{n}+\frac{i}{\omega_{\alpha}}U_{\alpha n}\,\hat{p}_{n}\right)\leavevmode\nobreak\ ,$
$\hat{a}_{\alpha}^{\dagger}$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sqrt{\frac{\omega_{\alpha}}{2}}\leavevmode\nobreak\ \sum_{n}\left(\hat{q}_{n}\left(U^{\dagger}\right)_{n\alpha}-\frac{i}{\omega_{\alpha}}\hat{p}_{n}\left(U^{\dagger}\right)_{n\alpha}\right)\leavevmode\nobreak\ .$ (74)
@@
$\varphi(x_{n})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \hat{q}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\sqrt{2\omega}}(\hat{a}_{n}+\hat{a}_{n}^{\dagger})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}}}\left(U_{n\alpha}^{\dagger}\hat{a}_{\alpha}+\hat{a}^{\dagger}_{\alpha}U_{\alpha n}\right)\leavevmode\nobreak\ ,$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-i\alpha n}+\hat{a}^{\dagger}_{l}e^{i\alpha n}\right)\leavevmode\nobreak\ ,$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}+\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\leavevmode\nobreak\ ,$ (75)
@@
$\pi(x_{n})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \hat{p}_{n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{\alpha}(-i)\sqrt{\frac{\omega_{\alpha}}{2aN}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}-\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\leavevmode\nobreak\ .$ (76)
[…]
$\varphi(x)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \int\frac{\textrm{d}k}{(2\pi)}\frac{1}{\sqrt{2\omega_{k}}}\left(\hat{a}_{k}e^{ikx}+\hat{a}^{\dagger}_{k}e^{-ikx}\right)\leavevmode\nobreak\ ,$
$\pi(x)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \int\frac{\textrm{d}k}{(2\pi)}(-i)\sqrt{\frac{\omega_{k}}{2}}\left(\hat{a}_{k}e^{ikx}-\hat{a}^{\dagger}_{k}e^{-ikx}\right)\leavevmode\nobreak\ ,$ (77)
where $\omega_{k}=\sqrt{\omega^{2}+k^{2}}$, and where we have made the usual rescaling of the ladder operators by the square-root of the volume, namely $\sqrt{a}\,\hat{a}_{\alpha}\longrightarrow\hat{a}_{k}/{\sqrt{aN}}$, in order to be consistent with $[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}]=2\pi a\delta(k-k^{\prime})$, versus $[\hat{a}_{\alpha},\hat{a}_{\beta}^{\dagger}]=\delta_{\alpha\beta}$.

### v2

@@ IV.2.2 Phase transitions
Figure 16: Phase transitions on the qumode lattice for two different parameter settings showing the measured expectation value $\langle\varphi\rangle$.
@@
Now let us turn to phase transitions in the system as a whole. In Fig. 16 we show two examples of phase transitions that were produced with the lattice initialised in the random sampling approximation to the QFT false vacuum that was discussed in Sec. III.1. The potential used for these examples is the $\tanh q^{\ell}$ function of Eq. (24) with $\ell=2$. The true minimum for these examples was chosen to be at $\varphi_{\rm tv}=0.5$ in order to allow for relatively rapid transitions. The figures show the production of bubbles which coalesce to complete the phase transition. The first example with $\varepsilon=0.5$ is very rapid because the instanton is very small (with $r_{c}\approx 2$). The second example has $\varepsilon=0.4$ and the instanton size has grown enough ($r_{c}\approx 5$) to noticeably slow the bubble production, as would be anticipated from the analysis in Sec. IV.1. One observation that supports the bubble formation being the result of genuine instanton transitions is that we find that the maximum bond dimension has to exceed a certain threshold in order for this behaviour to appear (if the dimension is set to be too small then there is little tunnelling). Moreover, this threshold appears to grow with the instanton size. In future work, an in-depth study of these relations will be carried out, with a particular focus on confirming the exponential dependence of the decay rate $\Gamma$ on the action.
IV.3 Seeding an instanton in the full QFT
Finally, we present a tunnelling process in the genuine QFT metastable vacuum initialised by imaginary-time TEBD as discussed in Sec. III.2. As a test we wish to seed the production of an instanton using a small perturbation. To do this we first prepare the initial state as per Sec. III.2 in the metastable QFT vacuum around $\varphi_{\rm fv}$. We then apply a small seed to stimulate the nucleation of a bubble within a reasonable time. For direct comparison we will consider bubble nucleation in the $\tanh q^{\ell}$ potential of Eq. (24) with $\lambda=\mu=1$, $\ell=2$ and parameters $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.3$. This is the same potential that was used to produce the ideal false-vacuum decay plot in Fig. 13.
In order to find a suitable seed without just randomly stimulating classical (equivalently thermal) transition, we must determine the optimal direction in field-space in which to push the field. We can do this by considering the expansion of the field around the $O(2)$ bounce solution described in Sec. IV.1. In particular we would like to identify a negative eigenmode whose instability leads it to grow exponentially, thereby driving the decay of the false vacuum. We will refer to this as the “negative mode”.
There is a well-known procedure for this, which is to linearise the fluctuations around the bounce solution, $\varphi_{B}(x)$, as
$\varphi(x,t)\penalty 10000\ =\penalty 10000\ \varphi_{B}(x,t)+\eta(x,t)\penalty 10000\ ,$ (76)
such that there is an effective $x$-dependent Lagrangian for $\eta$ of the form
${\cal L}\penalty 10000\ =\penalty 10000\ \frac{\eta_{t}^{2}}{2}-\frac{{\eta}_{x}^{2}}{2}-\frac{V^{\prime\prime}(\varphi_{B}(x))}{2}\eta^{2}\penalty 10000\ .$ (77)
Separating variables, $\eta=\chi(x)\tau(t)$, the equations of motion give $\tau_{tt}=\omega^{2}_{-}\tau$, and
$-\chi_{xx}\penalty 10000\ +\penalty 10000\ (V^{\prime\prime}(\varphi_{B}(x))+\omega_{-}^{2})\chi\penalty 10000\ =\penalty 10000\ 0\penalty 10000\ ,$ (78)
where we choose the separation constant $\omega_{-}$ anticipating exponential growth of the form $\tau(t)=Ae^{\omega_{-}t}+Be^{-\omega_{-}t}$. Clearly this can be consistent only if the mode in question $\chi(x)$ is a bound state in $x$ (as opposed to say an extended oscillatory mode in $x$). As Eq. (78) is simply the time-independent Schödinger equation with an effective potential $V_{\rm eff}(x)=V^{\prime\prime}(\varphi_{B}(x))$ and energy $\omega_{-}^{2}$, this is only possible if there are potential wells associated with regions of negative $V_{\rm eff}$. Thus a single exponentially growing mode, namely the negative mode, will be trapped in the regions with $V^{\prime\prime}(\varphi_{B})<0$, and indeed solving Eq. (78) will determine the “ground state energy” and hence the value of $\omega_{-}$: these regions correspond to those parts of $\varphi_{B}(x)$ that are beneath the potential barrier, which are in turn precisely the locations of the bubble walls of the bounce solution. We show this negative mode in Fig. 17
Figure 17: The negative mode for false vacuum decay in the $\tanh q^{\ell}$ potential of Eq. (24) with $\lambda=\mu=1$, $\ell=2$ and parameters $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.3$. The mode is trapped in the “effective potential wells” associated with the walls of the $O(2)$ bounce solution.
Generally, the negative mode will grow as
$\eta(x,t)\penalty 10000\ =\penalty 10000\ \left(q_{0}\cosh(\omega_{-}t)+\frac{p_{0}}{\omega_{-}}\sinh(\omega_{-}t)\right)\penalty 10000\ \chi(x)\penalty 10000\ ,$ (79)
with $q_{0}$ and $p_{0}$ indicating displacement and momentum at $t=0$. Thus the desired seed consists of a small kick to the false vacuum in the direction of this mode. A suitable and straightforward implementation then is to apply a momentum kick to our vacuum MPS of the form
$\hat{U}_{p_{0}}\penalty 10000\ =\penalty 10000\ e^{ip_{0}\chi(x_{i})\hat{q}_{i}}\penalty 10000\ ,$ (80)
where $\chi(x)$ is understood to be the normalised relevant bound state solution to Eq. (78), and where $p_{0}$ is a constant, which in our studies we take to be $p_{0}=0.1$.
The subsequent evolution of the scalar field VEV is shown in Fig. 18, where we observe behaviour that closely follows the theoretical prediction of Sec. IV.1 and the ideal case (with identical parameters) in Fig. 13. As expected there is an initial period during which nothing appears to happen, in which the negative mode is growing exponentially. Given the behaviour of this mode in Eq. (79), the time for instability to set in is given by $1/\omega_{-}$ with only logarithmic dependence on the initial kick size $p_{0}$, provided that it is large enough. (As we are not actually in the background of the critical bubble, an insufficiently large $p_{0}$ just perturbs the system in the harmonic basin of the false vacuum, with the energy radiating away.) As anticipated the bubble appears “walls first”, with the interior being pulled to the true vacuum by the walls.
There are other features that we should remark on. First we note the residual negative mode which overshoots the saddle and ultimately disappears in radiation. We can also see a residual long-lived “oscillon” at the site of the decay [69, 70].
It is worth understanding the differences between the Fig. 13 and Fig. 18 as well. The qumode lattice simulation has a somewhat less well defined hyperbolic trajectory. This is due to the discretisation of the system onto a lattice. This causes the non-ideal dispersion relation in Eq. (95), which results in a softening of short wavelength modes. (In fact we can observe the same effect when we latticise the classical PDE solution.) A second difference is the slightly enhanced oscillations in Fig. 18: this is due to the seed that we introduced to stimulate the transition.
Figure 18: An isolated tunnelling event from the metastable QFT vacuum. The system is initialised in the metastable QFT vacuum with the $\tanh q^{\ell}$ potential of Eq. (24) with $\lambda=\mu=1$, $\ell=2$ and parameters $\varphi_{\rm fv}=-2$, $\varphi_{\rm tv}=0.7$, $\varepsilon=0.3$. A small seed is applied (consisting of a small momentum boost applied at $t=0$ along the exponentially growing “negative mode” in Fig. 17) in order to stimulate an instantonic transition. Here $\langle\varphi\rangle$ is the measured expectation value.
V Summary and Conclusions
@@
In this work, we have presented a novel framework for simulating false vacuum decay in scalar QFT using a qumode-based Hamiltonian formulation. Our method overcomes longstanding challenges associated with simulating highly entangled, non-equilibrium tunnelling dynamics in QFT by leveraging a qumode lattice representation and implementing real-time evolution via tensor network techniques, specifically the time-evolving block decimation algorithm.
@@
Unlike previous approaches, which typically restricted attention to spin chains or Euclidean simulations, our framework, based on Hamiltonian simulation, remains firmly within the QFT setting and directly simulates real-time tunnelling phenomena. The core innovation lies in the use of qumodes, that is bosonic modes with (quasi-)continuous Hilbert spaces, which enable the construction of scalar field theories with arbitrary potentials. This setup can, in principle, be implemented on continuous-variable quantum computing devices in its presented form. Simulation on current classical hardware may be performed by truncating the local Hilbert space and employing a matrix product state ansatz for the quantum state. Such truncations would not be needed on continuous-variable quantum hardware, which would enable simulations of genuine Coleman tunnelling.
This work made two key advances. The first was an algorithm for preparing the full initial state vacuum of even strongly coupled QFTs using imaginary-time TEBD, thus enabling scalable simulations with tunable entanglement control. The second key advance was the identification and study of the role of the negative tachyonic mode in false vacuum decay. This goes well beyond what is possible in the traditional treatments of false vacuum decay, exposing the intrinsically quantum origin of the non-perturbative instability.
@@
We demonstrated the validity of our method through a sequence of other complex tests. Importantly, we showed that quantum bubble nucleation emerges in this framework only once sufficient entanglement is captured by the MPS. The bond dimension required increases with bubble size, suggesting a direct numerical signature of the entanglement entropy scaling expected for false vacuum decay. This provides a powerful diagnostic and highlights the necessity of entanglement-aware methods in simulating quantum phase transitions.
@@
Compared to previous literature, our work represents the first demonstration of scalar QFT vacuum decay using a real-time evolved Hamiltonian lattice system with quasi-continuous local degrees of freedom. While continuous-variable tensor network methods, such as continuous MPS or cPEPS [85, 86], have been developed for ground-state field theory, and continuous-variable quantum computing proposals exist for QFT [25, 87], no prior work has combined qumode-lattice discretisation with real-time tensor-network evolution to simulate false vacuum decay in QFT. Our method is thus, to the best of our knowledge, the first demonstration of qumode tensor networks applied to non-equilibrium tunnelling dynamics in scalar field theory.
@@
It can open the door to systematic studies of vacuum decay in arbitrary scalar potentials, including non-renormalisable theories and those beyond the reach of semiclassical methods. Future extensions include the exploration of multi-field tunnelling landscapes, and the application of this method to higher-dimensional QFTs. Our results show the viability of qumode tensor networks as a bridge between classical simulation and future quantum computing applications in non-perturbative quantum field theory.
@@ A.1 Free fields on the qumode lattice
$\mathscr{L}\penalty 10000\ =\penalty 10000\ \frac{1}{2}\left(\partial^{\mu}\varphi(x,t)\partial_{\mu}\varphi(x,t)-\omega^{2}\varphi(x,t)^{2}\right)\penalty 10000\ ,$ (81)
@@
$\pi(x,t)\penalty 10000\ =\penalty 10000\ \frac{\partial\mathscr{L}}{\partial(\partial_{0}\varphi(x,t))}\penalty 10000\ =\penalty 10000\ \partial_{0}\varphi(x,t)\penalty 10000\ ,$ (82)
@@
$\mathscr{H}(x,t)\penalty 10000\ =\penalty 10000\ \frac{1}{2}\left(\pi(x,t)^{2}+(\nabla\varphi(x,t))^{2}+\omega^{2}\varphi(x,t)^{2}\right)\penalty 10000\ .$ (83)
@@
$\hat{q}_{n}(t)\penalty 10000\ \leftrightarrow\penalty 10000\ \varphi(x_{n},t)\penalty 10000\ ,\qquad\hat{p}_{n}(t)\penalty 10000\ \leftrightarrow\penalty 10000\ \pi(x_{n},t)\penalty 10000\ .$ (84)
@@
$\partial_{i}\varphi(x_{n},t)\penalty 10000\ \leftrightarrow\penalty 10000\ \frac{{\hat{q}}_{n+1}(t)-{\hat{q}}_{n}(t)}{a}\penalty 10000\ .$ (85)
@@
$H\penalty 10000\ =\penalty 10000\ \frac{a}{2}\sum_{n=0}^{N-1}\left[{\hat{p}}_{n}(t)^{2}+\left(\frac{\hat{q}_{\,\overline{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}+\omega^{2}\hat{q}_{n}(t)^{2}\right]\penalty 10000\ ,$ (86)
@@
$\hat{q}_{n},\,\hat{p}_{m}]\penalty 10000\ =\penalty 10000\ i\delta_{nm}\penalty 10000\ ,$ (87)
@@
$\overline{n+1}\penalty 10000\ =\penalty 10000\ n+1\mod(N)\penalty 10000\ .$ (88)
@@
$|\Psi\rangle\penalty 10000\ =\penalty 10000\ \otimes_{n}|\psi_{n}\rangle\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ |{\mathbf{q}}\rangle\penalty 10000\ =\penalty 10000\ \otimes_{n}|q_{n}\rangle\penalty 10000\ .$ (89)
@@
$\langle{\mathbf{q}}|\left(\frac{\hat{q}_{\,\overline{n+1}}(t)-\hat{q}_{n}(t)}{a}\right)^{2}|\Psi_{\rm degen}\rangle\penalty 10000\ =\penalty 10000\ 0\penalty 10000\ .$ (90)
@@
$\langle{\mathbf{q}}|\Psi_{\rm degen}\rangle\penalty 10000\ =\penalty 10000\ \prod_{n}\langle q_{n}|\psi_{n}\rangle\penalty 10000\ .$ (91)
@@
$U_{\alpha m}\penalty 10000\ =\penalty 10000\ \frac{1}{\sqrt{N}}\nu^{\alpha m}\penalty 10000\ ,$ (92)
@@
$H\penalty 10000\ =\penalty 10000\ \frac{a}{2}\sum^{N-1}_{\alpha=0}\left(|\hat{p}_{\alpha}|^{2}+\omega_{\alpha}^{2}|\hat{q}_{\alpha}|^{2}\right)\penalty 10000\ ,\\$ (93)
@@
$\hat{p}_{\alpha}\penalty 10000\ =\penalty 10000\ \sum_{n}U_{\alpha n}\hat{p}_{n}\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{q}_{\alpha}\penalty 10000\ =\penalty 10000\ \sum_{n}U_{\alpha n}\hat{q}_{n}\penalty 10000\ ,$ (94)
@@
$\omega_{\alpha}\penalty 10000\ =\penalty 10000\ \sqrt{\left(\omega^{2}+\frac{4}{a^{2}}\sin^{2}\left(\frac{\pi\alpha}{N}\right)\right)}\penalty 10000\ .$ (95)
@@
$k_{\alpha}\penalty 10000\ =\penalty 10000\ 2\pi\frac{\alpha}{aN}\penalty 10000\ =\penalty 10000\ 2\pi\frac{\alpha}{L}\penalty 10000\ ,$ (96)
@@
$\hat{a}_{n}$ $\penalty 10000\ =\penalty 10000\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}+\frac{i}{\omega}\hat{p}_{n}\right)\penalty 10000\ ,$
$\hat{a}^{\dagger}_{n}$ $\penalty 10000\ =\penalty 10000\ \sqrt{\frac{\omega}{2}}\left(\hat{q}_{n}-\frac{i}{\omega}\hat{p}_{n}\right)\penalty 10000\ ,$ (97)
@@
$\hat{a}_{\alpha}$ $\penalty 10000\ =\penalty 10000\ \sqrt{\frac{\omega_{\alpha}}{2}}\penalty 10000\ \sum_{n}\left(U_{\alpha n}\,\hat{q}_{n}+\frac{i}{\omega_{\alpha}}U_{\alpha n}\,\hat{p}_{n}\right)\penalty 10000\ ,$
$\hat{a}_{\alpha}^{\dagger}$ $\penalty 10000\ =\penalty 10000\ \sqrt{\frac{\omega_{\alpha}}{2}}\penalty 10000\ \sum_{n}\left(\hat{q}_{n}\left(U^{\dagger}\right)_{n\alpha}-\frac{i}{\omega_{\alpha}}\hat{p}_{n}\left(U^{\dagger}\right)_{n\alpha}\right)\penalty 10000\ .$ (98)
@@
$\varphi(x_{n})$ $\penalty 10000\ =\penalty 10000\ \hat{q}_{n}\penalty 10000\ =\penalty 10000\ \frac{1}{\sqrt{2\omega}}(\hat{a}_{n}+\hat{a}_{n}^{\dagger})$
$\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}}}\left(U_{n\alpha}^{\dagger}\hat{a}_{\alpha}+\hat{a}^{\dagger}_{\alpha}U_{\alpha n}\right)\penalty 10000\ ,$
$\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-i\alpha n}+\hat{a}^{\dagger}_{l}e^{i\alpha n}\right)\penalty 10000\ ,$
$\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{\sqrt{2\omega_{\alpha}\,N}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}+\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\penalty 10000\ ,$ (99)
@@
$\pi(x_{n})$ $\penalty 10000\ =\penalty 10000\ \hat{p}_{n}\penalty 10000\ =\penalty 10000\ \sum_{\alpha}(-i)\sqrt{\frac{\omega_{\alpha}}{2aN}}\left(\hat{a}_{\alpha}e^{-ik_{\alpha}x_{n}}-\hat{a}^{\dagger}_{\alpha}e^{ik_{\alpha}x_{n}}\right)\penalty 10000\ .$ (100)
Taking the continuum limit as $N\rightarrow\infty$, and replacing $\sum_{\alpha}\equiv\frac{aN}{2\pi}\int dk$ we retrieve the standard QFT definition of the field as an expansion in momentum space at a fixed time, namely
$\varphi(x)$ $\penalty 10000\ =\penalty 10000\ \int\frac{\textrm{d}k}{(2\pi)}\frac{1}{\sqrt{2\omega_{k}}}\left(\hat{a}_{k}e^{ikx}+\hat{a}^{\dagger}_{k}e^{-ikx}\right)\penalty 10000\ ,$
$\pi(x)$ $\penalty 10000\ =\penalty 10000\ \int\frac{\textrm{d}k}{(2\pi)}(-i)\sqrt{\frac{\omega_{k}}{2}}\left(\hat{a}_{k}e^{ikx}-\hat{a}^{\dagger}_{k}e^{-ikx}\right)\penalty 10000\ ,$ (101)
where $\omega_{k}=\sqrt{\omega^{2}+k^{2}}$, and where we have made the usual rescaling of the ladder operators by the square-root of the volume, namely $\sqrt{a}\,\hat{a}_{\alpha}\longrightarrow\hat{a}_{k}/{\sqrt{aN}}$, in order to be consistent with $\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}]=2\pi a\delta(k-k^{\prime})$, versus $\hat{a}_{\alpha},\hat{a}_{\beta}^{\dagger}]=\delta_{\alpha\beta}$.